# suppSupplementary Material References

### v1

@@ 4 Research Design
Our analyses are based on a dataset covering 120 autocratic countries over the period 1951–2018 from Choulis, Escribà-Folch, and Mehrl (2024). The target variable $y_{i,t}$ is a binary indicator of secret police onset in country $i$ and year $t$, taking the value 1 if a secret police is established and 0 else. Following McGrath (2015), we set $y_{i,t}$ to missing if a secret police exists but was established before. If the transition is interrupted by a gap in the dataset due to the suspension of autocracy, this is not counted as a formation. As a result, we disregard any formations at time point $t$, if, e.g., there was no autocracy or observation at $t-1$. According to this scheme, we observe 31 secret police formations in 29 countries, with Iran and Cuba each experiencing two formations. The dataset contains 3,759 observations in total. We include all independent variables listed in Table 1 lagged by one year to account for the time between changes in the covariates and their potential effect on secret police formation. Missing values are imputed using bootstrap-based methods with the Amelia package (Honaker et al., 2011) [^3] For this imputation step, we require that the missing values are missing at random (MAR), meaning that the probability of missingness depends only on observed data and not on the missing values themselves. This is a plausible assumption for our data.. Thus, we make use of multiple imputation in an iterative algorithm that, first, imputes missing values and, second, maximizes the likelihood of the complete data. To account for uncertainty introduced the imputation step, we adopt a bootstrapping approach. For robustness, we impute the dataset five times, apply our methodology to each imputed version, and then average the resulting models to obtain more stable estimates.
@@
The employed methodology follows a two-step approach: first, we identify the most relevant covariates via statistical variable selection, and second, we examine the magnitude, direction, and significance of their effects. The second step simply amounts to estimating logistic regression models, as we fit a logistic model on the selected variables for each dataset and pool the results by averaging. However, the first step requires a more detailed explanantion.
@@
Since the sample size of our data is relatively small and our aim is to identify a parsimonious set of covariates from Table 1, we first fit a Lasso model to the data where the absolute value of all parameters is penalized with a penalization parameter $\lambda$ (Tibshirani, 1996). This penalization will set the parameters of some coefficients to zero and thus detects which variables are associated with the establishment of secret police organizations. Since the dependent variable is binary, we employ a logit link function. In line with standard practice (Tibshirani, 1996; Hastie et al., 2009; Friedman et al., 2010), the penalization parameter $\lambda$ was chosen by minimizing the deviance through three-fold cross validation. Given that there are only 31 observed formations, we stratify each cross-validation dataset to ensure that each fold contains approximately the same number of secret police establishments. This ensures a balanced distribution across the folds. The Lasso model is applied to each of the five imputed datasets detailed in the previous paragraph. We keep only those variables that appear in at least three out of the five models fitted on the different imputed datasets. Finally, we fit an unconstrained logistic model with the selected variables for each dataset and pool the results via Rubin’s (1987) rule.
@@
To further validate our findings, we repeat the analysis using an alternative stepwise selection approach based on Akaike Information Criterion (AIC) instead of Lasso. This procedure begins with an empty model and iteratively adds the variable that maximizes the AIC at each step until no further improvement is possible. Although the stepwise approach is straightforward and easy to interpret, it makes local, step-by-step decisions that may not yield the overall optimal set of predictors. Moreover, its results can be sensitive to the order in which variables are introduced (Hastie et al., 2020). In contrast, the Lasso method is more stable with strong theoretical foundations. Therefore, we use the stepwise approach only to validate the model. The variables selected via the stepwise approach closely align with those identified by the Lasso method, as shown in Table 1.
@@
As a robustness check, we replace the logit-link with a complementary log-log (cloglog) link function. Unlike the symmetric logit transformation, the cloglog-link is asymmetric and commonly used for modeling binary outcomes where one class is rare (Tutz and Schmid, 2016), as is the case with our dependent variable. The results using the cloglog-link remain similar to those from models fitted with the logit-link.
@@ 5 Empirical Results
The first set of results, that is, which variables were selected by which approach (Lasso with logit link, Lasso with cloglog link, stepwise selection with logit link, and stepwise selection with cloglog link models), is provided in Table 1. We observe that several of the potential predictor variables were not selected by all four selection approaches. None of the several measures of state capacity was selected in a single model, suggesting that this institutional dimension is not associated with autocrats’ decision to set up secret police. There is also no evidence that the occurrence or a history of coup attempts or civil war influence whether secret police are created. Further, existing alternative approaches to managing threats to the regime, such as counterbalancing or clientelism, were left out of all models, as were demographic variables. Finally, regional protests and coup attempts do not appear to play a role in creating secret police. At the same time, Table 1 exhibits largely compatible results between the four variable selection approaches, indicating that overall results do not depend on a particular selection approach or link function.
@@
Figure 1 displays the marginal effects on the probability of secret police formation for all variables selected by at least one of the employed methods (Leeper, 2024) [^4] For full results tables, see the Appendix.. We restrict the plot to variables identified by at least one of the four approaches detailed in Table 1. A first takeaway from Figure 1 is that none of the detected effect sizes is large – however, this is unsurprising given how rare secret police formations are.
@@
Figure 1: Predictors of Secret Police:Marginal effect estimates are shown with 90% and 95% confidence intervals, represented by thick and thin whiskers, respectively. Each variable is associated with four coefficient estimates, corresponding to LASSO and Stepwise selection methods applied with logit and cloglog link functions. Coefficients for variables not selected by a given method are plotted as zero to maintain comparability across approaches.
@@
Looking at specific effect estimates, Figure 1 supports several of the intuitions developed in Section 3 about secret police formation. First, the autocrat’s threat environment appears to be important: secret police are more likely to be established if protests are more rife, civil society organisations challenging the system of government exist and are stronger, and if the country is operating within an international rivalry. Interestingly, this positive relationship between threat and secret police formation appears limited to more latent, not yet escalated threats. This interpretation is supported by the finding that variables capturing intrastate conflict, coup attempts, and the presence of interstate dispute remain unselected, while the count of interstate disputes is selected only in the two stepwise selection models where it is found to have a negative relationship with secret police being established. This conclusion may suggest that secret police are formed to pre-empt and be prepared for such escalation when other threats make it appear likely, as opposed to when it actually takes place. In other words, secret police may be created to delay or even stop an endgame from occurring, as opposed to helping them survive the endgame when it does occur. Indeed, when faced with a potential endgame, creating new coercive institutions may only serve to further weaken the ruler’s position vis-a-vis the ruling elite instead of strengthening it (see Svolik, 2012), or be unattractive as the ruler actually enjoys increased elite loyalty for the time being (McMahon and Slantchev, 2015). And finally, the result that none of the several indicators of coup attempts and coup risk appear to matter, while protests and anti-system civil society organizations do, may suggest that secret police are created more to target threats emanating from the broader citizenry, as opposed to (military) elites.
Moreover, the results in Figure 1 indicate that secret police formation is not associated with a country’s respect for physical human rights and is less likely if, specifically, civil society is already being repressed. Looking at regime characteristics, the results presented in Figure 1 highlight that secret police are increasingly less likely to be established as a regime is in place for longer – suggesting that once rule has been truly established over time, a secret police either already exists or is not necessary anymore as the regime is protected via other means. Along these lines, there is also some evidence, though found only in the stepwise selection models, that autocrats are more likely to establish secret police as they concentrate more power on themselves, mirroring similar findings for other kinds of personalised security institutions (e.g. Escribà-Folch et al., 2020), but less likely if they can rely on ideological claims to legitimize their rule.
While none of the state capacity indicators were selected, there is support for the idea that the regime’s material resources affect the establishment of a secret police. Military spending is positively associated with secret police formation, as is GDP per capita. Gas production, however, has a negative coefficient estimate, where selected. Unsurprisingly, wealthier regimes and those that already spend more on security are accordingly also more likely to invest in creating a secret police. Finally, the variable capturing the wider participation level in civil society was selected in two models, but its coefficient estimate allows for little interpretation.
In sum, our empirical results thus suggest that autocrats establish secret police when facing a particular threat environment characterized by domestic opposition from its citizens, as well as the latent external dangers captured by international rivalry. In contrast, other treats, including coups and civil war, appear not to matter, suggesting that secret police may be created to pre-empt, rather than tackle actualized threats, and to target threats from ciizens rather than elites. Additionally, our results highlight that autocrats are more likely to establish secret police early in their tenure, when they have sufficient financial resources, and when they spend those resources on security. As such, our results provide novel insights into why, as highlighted above, secret police might be so surprisingly rare in autocracies: Rulers have to face a fairly specific combination of threats and their own capabilities in terms of resourcing to be both able and willing to create a secret police. In particular, secret police are more likely to be formed in response to mass-based threats that have not (yet) escalated into civil war, but also when the ruler faces external, un-escalated threats in the form of rivalries. But to establish such a security institution, the ruler also requires sufficient material resourcing and, at least according to some of our models, sufficient power vis-a-vis their ruling coalition, thus also pointing back to studies highlighting the internal dynamics of secret police institutions and the risks associated with their existence (Egorov and Sonin, 2011; Zakharov, 2016; Dragu and Przeworski, 2019).
@@ 6 Conclusion
Secret police play a key role in ensuring authoritarian rulers’ survival. Nonetheless, secret police exist in not even a third of non-democratic country-years where data is available to us. Motivated by this apparent disconnect between the usefulness and commonness of secret police, we investigate when secret police are established in autocracies. Therefore, we collected a list of potential driving factors from the growing literature on state security force structure, added several further variables motivated by insights from broader research on authoritarian survival, and used variable selection methods to comprehend which included covariates are associated with secret police formation. Our results highlight that secret police are most likely to be created when rulers face particular, not yet escalated threats motivating their establishment but also have the material resources to do so. In particular, secret police appear to be established in response to mass- and civil society-based dissent, but not to a history of coup attempts or civil war. And looking at external threats, international rivalries are consistently found to increase the probability of secret police formation whereas actual international conflict is not.
These results contribute to our understanding of secret police, state security force structure, and autocrats’ tools of survival in several ways: We challenge the often-held assumption that secret police are simply a static feature of autocracies, highlighting instead that they are absent more often than not. Further, we provide both a set of empirical correlates and an overarching intuition to understand why this is the case. As such, this research not only highlights the need for more theory-driven work on when secret police are established but also provides a strong empirical foundation for it. Finally, it highlights that not only autocrats’ choices regarding their political institutions (e.g. Meng et al., 2023) but also their security apparatus require study, as the latter exhibits both variation and clear empirical consequences in terms of repression and regime survival.
Appendix
Table A.1: Full results table underlying figure 1. Coefficients for covariates that were not selected by the Lasso model are represented by the symbol —.
LASSO Stepwise
  Variable Logit Cloglog Logit Cloglog
Intercept --7.7292 --7.6051 --10.8003 --10.5246

### v2

@@ 4 Research Design
Our analyses are based on a dataset covering 120 autocratic countries over the period 1951–2018 from Choulis, Escribà-Folch, and Mehrl (2024). The target variable $y_{i,t}$ is a binary indicator of secret police onset in country $i$ and year $t$, taking the value 1 if a secret police is established and 0 otherwise, where secret police are coded in line with the definition introduced above along five criteria capturing their targeting of political opponents, operational independence and secrecy, specialisation in intelligence operations, and regular use of violent policing practices (Choulis et al., 2024; Mehrl and Choulis, 2024). Following McGrath (2015), we set $y_{i,t}$ to missing if a secret police exists but was established before. If the transition is interrupted by a gap in the dataset due to the suspension of autocracy, this is not counted as a formation. As a result, we disregard any formations at time point $t$, if, e.g., there was no autocracy or observation at t1t-1. According to this scheme, we observe 31 secret police formations in 29 countries, with Iran and Cuba each experiencing two formations. The dataset contains 3,759 observations in total. We include all independent variables listed in Table 2 lagged by one year to account for the time between changes in the covariates and their potential effect on secret police formation. Missing values are imputed using bootstrap-based methods with the Amelia package (Honaker et al., 2011) [^5] For this imputation step, we require that the missing values are missing at random (MAR), meaning that the probability of missingness depends only on observed data and not on the missing values themselves. This is a plausible assumption for our data.. Thus, we make use of multiple imputation in an iterative algorithm that, first, imputes missing values and, second, maximises the likelihood of the complete data. To account for uncertainty introduced the imputation step, we adopt a bootstrapping approach. For robustness, we impute the dataset five times, apply our methodology to each imputed version, and then average the resulting models to obtain more stable estimates. As we are interested in when and where secret police are formed, we make use of both variation across countries and across time.
@@
The employed methodology follows a two-step approach: first, we identify the most relevant covariates via statistical variable selection, and second, we examine the magnitude, direction, and significance of their effects. The second step simply amounts to estimating logistic regression models, as we fit a logistic model on the selected variables for each dataset and pool the results by averaging. However, the first step requires a more detailed explanation.
@@
Since the sample size of our data is relatively small and our aim is to identify a parsimonious set of covariates from Table 2, we first fit a Lasso model to the data where the absolute value of all parameters is penalised with a parameter $\lambda$ (Tibshirani, 1996). This penalisation will set the parameters of some coefficients to zero and thus detects which variables are associated with the establishment of secret police organisations. Since the dependent variable is binary, we employ a logit link function. In line with standard practice (Tibshirani, 1996; Hastie et al., 2009; Friedman et al., 2010), the penalisation parameter $\lambda$ was chosen by minimising the deviance through three-fold cross validation. Given that there are only 31 observed formations, we modify the standard Lasso by stratifying each cross-validation dataset to ensure that every fold contains approximately the same number of secret police establishments. This ensures a balanced distribution across the folds, and makes the Lasso model also applicable to binary outcomes where one value is rare. The stratified Lasso model is applied to each of the five imputed datasets detailed in the previous paragraph. We keep only those variables that appear in at least three out of the five models fitted on the different imputed datasets. Finally, we fit an unconstrained logistic model with the selected variables for each dataset and pool the results via Rubin’s (1987) rule.
@@
To further validate our findings, we repeat the analysis using an alternative stepwise selection approach based on Akaike Information Criterion (AIC) instead of Lasso. This procedure begins with an empty model and iteratively adds the variable minimising the AIC at each step until no further improvement is possible. While straightforward and easy to interpret, this approach makes local, step-by-step decisions that may not yield the overall optimal set of predictors. Moreover, results are often sensitive to the order in which variables are introduced (Hastie et al., 2020). As such, we prioritise the Lasso and use the stepwise selection approach as a robustness check, ensuring that our results are not specific to one particular variable selection method. Reassuringly, Table 2 shows that the variables selected via the two approaches are closely aligned.
@@
As a further robustness check, we replace the logit-link with a complementary log-log (cloglog) link function. Unlike the symmetric logit transformation, the cloglog-link is asymmetric and commonly used for modelling binary outcomes where one class is rare (Tutz and Schmid, 2016), as is the case with our dependent variable. The results using the cloglog-link remain similar to those from models fitted with the logit-link.
@@ 5 Empirical Results
The first set of results, that is, which variables were selected by which approach (Lasso with logit link, Lasso with cloglog link, stepwise selection with logit link, and stepwise selection with cloglog link models), is provided in Table 2. We observe that several of the potential predictor variables were not selected by all four selection approaches. None of the several measures of state capacity was selected in a single model, suggesting that this institutional dimension is not associated with autocrats’ decision to set up secret police. There is also no evidence that the occurrence or a history of coup attempts or civil war influence whether secret police are created. Further, existing alternative approaches to managing threats to the regime, such as counterbalancing or clientelism, were left out of all models, as were demographic variables. Finally, intrastate conflict and protests in the neighbourhood do not appear to play a role for creating secret police. At the same time, Table 2 exhibits largely compatible results between the four variable selection approaches, indicating that overall results do not depend on a particular selection approach or link function.
@@
Figure 1 displays the marginal effects on the probability of secret police formation for all variables selected by at least one of the employed methods [^6] See the Supplementary Materials for results tables.. We restrict the plot to variables identified by at least one of the four approaches detailed in Table 2. Figure 1 indicates that none of the detected effect sizes is large – however, this is unsurprising given how rare secret police formations are.
@@
Figure 1: Predictors of Secret Police: Marginal effect estimates with 90% and 95% confidence intervals (thick and thin whiskers, respectively). Each variable is associated with four coefficient estimates, corresponding to LASSO and Stepwise selection methods applied with logit and cloglog link functions. Coefficients for unselected variables plotted as zero to maintain comparability across approaches.
@@
Looking at specific effect estimates, Figure 1 supports several of the intuitions developed in Section 3 about secret police formation. First, the autocrat’s external threat environment appears to be important: secret police are more likely to be established if civil society organisations challenging the system of government exist and are stronger, protests are more rife, and if the country is operating within an international rivalry. Interestingly, this positive relationship between threat and secret police formation appears limited to more latent, structural threats which have not yet fully escalated. This interpretation is supported by the finding that variables capturing more shock-like events, such as intrastate conflict, coup attempts, and the presence of interstate dispute, remain unselected, while the count of interstate disputes is found to have a negative relationship with secret police establishment. The exception to this pattern is the positive coefficient estimate recovered for successful coups in the neighbourhood, which, though shock-like, indicates that autocrats observe such events, update their beliefs about the threats they are facing themselves, and move to adjust their security apparatus accordingly (Ben Hammou and Powell, 2025). In this sense, shock-like events elsewhere change the dictator’s belief about their own structural sources of danger. That successful coups in the neighbourhood are found to increase the probability of secret police establishment, while none of the domestic indicators of coup risk are selected, may further imply that dictators form secret police to stop military threats to their rule from emerging, rather than to tackle them once such threats are realised in the form of, for instance, politicised officer cliques within the military or rumors of coup plans. And more broadly, these conclusions suggest that secret police are formed to pre-empt and be prepared for escalation when the structural threat environment makes it appear likely, as opposed to when escalation actually takes place. In other words, secret police may be created to delay or even stop an endgame from occurring, not to help autocrats survive the endgame when it is already occurring.
Indeed, when faced with a potential endgame, creating new coercive institutions may only serve to further weaken the ruler’s position vis-a-vis the ruling elite instead of strengthening it (see Svolik, 2012), or be unattractive as the ruler actually enjoys increased elite loyalty for the time being (McMahon and Slantchev, 2015). This first set of results thus highlights the importance of the regime-external structural threat environment, which alters autocrats’ willingness and opportunity for policy changes, in predicting secret police formation. However, these threats appear to have to be realised to some extent, such as in the form of anti-system organisations or protests, as our results indicate structural predictors of threats, including recent intrastate conflict experience, civil society strength, or excluded ethnic population share, to be of negligible importance.
Regarding regime-internal structural characteristics, we find that secret police formation is less likely if levels of civil society repression are already high, and when the regime is in place for longer. As discussed above, both of these characteristics should reduce rulers’ perceived need, and hence their willingness, to invest in additional regime survival policies. In such cases, a secret police will thus either already exist or is not necessary any more as the regime is protected via other means. Along these lines, there is also some evidence, though found only select models, that autocrats are less likely to establish secret police if the specific leader has been in power for longer or they can rely on ideological claims to legitimise their rule. There is thus evidence that several regime-internal structural features, which we propose to reduce the need for and thus willingness to invest in secret police, to be associated with a lower probability of secret police formation. Interestingly, however, this finding does not extend to the presence of alternative security actors, such as paramilitaries, or sources of information and control, such as clientelism, as these remain unselected across models.
Moreover, while none of the state capacity indicators were selected, the results in Figure 1 offer support to the idea that the regime’s material resources act as a structural predictor of its ability to fund and maintain, and hence establish, a secret police. Specifically, military spending is positively associated with secret police formation, suggesting that as autocracies spend more on their security apparatus, they also have more opportunity to form a secret police. GDP per capita equally exhibits a positive effect estimate, suggesting a similar interpretation. Interestingly, oil and gas production remained generally unselected, indicating that it is not specifically disposable income which matters for whether secret police are established, but a regime’s general material resources. Finally, our models indicate that personalism, another structural characteristic of the regime which above is argued mainly affect their ruler’s ability to create secret police, matters: rulers who have concentrated more power on themselves are more likely to establish a secret police, with increasingly marginalised elites being less able to resist this move.
In the supplementary materials, we detail data sources and present a results table for the models underlying Figure 1. We apply these models to predict counterweight paramilitaries instead of secret police, and find that while many of the statistically significant variables are they same, their effect directions are flipped. This may indicate that rulers form these distinct security actors when facing different threats and conditions. And we estimate additional models where we additionally allow continuous covariates to enter the model as first differences and construct indicators of the time since their last realisation for all binary covariates.
Taken together, our empirical results suggest that autocrats establish secret police when facing a structural external threat environment, indicating opposition from their own citizens or international rivals. In contrast, other domestic threats, including coups and civil war, appear not to matter, suggesting that secret police may be created to pre-empt, rather than tackle actualised threats, and to mainly target threats from outside the regime rather than elites. Additionally, our results highlight that autocrats are more likely to establish secret police when their structural resource and power base allows them to do so, while the probability of secret police formation decreases if regime-internal structure indicates that such an additional security organisation is less needed.
@@ 6 Conclusion
Secret police, such as the Soviet Cheka, NKVD, and KGB, Franco’s Political-Social Brigade, and Saddam’s Mukhabarat, play a key role in ensuring authoritarian rulers’ survival. But despite this and the prominence of many such organisations, secret police exist in less than a third of non-democratic country-years where data is available to us. Motivated by this apparent disconnect between the usefulness and commonness of secret police, we investigate when secret police are established in autocracies. This paper develops a theoretical framework to identify and interpret predictors in a structured fashion, extends existing statistical variable selection techniques to better handle rare binary outcomes such as secret police formation, and applies the resulting method to assess which of the several potential predictors are actually meaningfully associated with the establishment of secret police.
Our results highlight that secret police are most likely to be formed when rulers face particular, structural rather than shock-like threats external to their regime, but also have the opportunity to do so. As such, our results point to why, as highlighted above, secret police might be so surprisingly rare in autocracies: Rulers have to face a fairly specific combination of threats and their own resources to be both able and willing to create a secret police. As such, secret police formation depends not only on rulers’ access to sufficient material resources to fund such an organisation, but also their exposure to threats which are regime-external and sufficiently important, but also not too acute. Otherwise, rulers appear to opt against creating secret police, potentially due to the very real costs these institutions can impose on rulers (Egorov and Sonin, 2011; Thomson, 2020). Connected to this, our data also suggest that an additional reason for secret police formation being so rare is that secret police tend to stick around. In other words, rulers have a hard time dissolving one and then establishing another secret police, further re-iterating that choosing to create such a powerful security agency is not an automatic decision for autocrats.
For future research, we offer both tools and directions to further advance our understanding of secret police, state security force structure, and autocrats’ strategies of survival. We provide a flexible theoretical framework to systematise and interpret potential predictor variables, thus allowing future work to straightforwardly add new covariates to those studied here. The stratified Lasso method developed for this research offers a directly applicable approach to variable selection for predicting secret police formation, but also a range of other relevant rare outcomes. And together with the data and code for this paper, these contributions will serve as a starting point for future studies that want to expand the range of considered predictor variables, investigate alternative dependent variables such as alternative security organisations or other institutional choices in autocracies, or more theoretically hone in on specific predictors with the aim of establishing causal relationships.
But additionally, given how scarcely secret police are formed and how enduring they are, this research also highlights their internal dynamics as a subject for future study. The recent wave of studies focusing on particular secret police cases in, for instance, Eastern Germany (Steinert, 2023), Poland (Hager and Krakowski, 2022; Thomson, 2020, 2024), Taiwan (Liu et al., 2026), and Argentina (Scharpf and Gläßel, 2020) is helpful in this regard. But future work should supplement this with a greater focus on comparison and, eventually, extending global data on secret police existence with more information regarding their activities, size, and surveillance methods.
Appendix A Data Sources and Software
Table A.1 lists the data sources for all independent variables included in the models presented here. As discussed in the main paper, the data on secret police existence used to construct the dependent variable ’secret police formation’ comes from Choulis, Escribà-Folch, and Mehrl (2024). In accessing some of the data sources listed in Table A.1, we benefited from Peacesciencer (Miller, 2022) and the WEP Dataverse (Graham and Tucker, 2019).
We utilised the packages cshapes (Schvitz et al., 2022) and sf (Pebesma and Bivand, 2023) for spatial processing. Multiple imputation of the missing values was facilitated by Amelia (Honaker et al., 2011), while glmnet was used for estimating Lasso regression and cross-validation. The margins Leeper (2024) package was used to calculate Average Marginal Effects (AME).
Table A.1: Variables and data sources
Variable Data source Variable Data source
Intrastate conflict: Dummy Davies et al. (2024) Clientelism Coppedge et al. (2023)
Intrastate conflict: Years since Davies et al. (2024) Ethnically excluded pop. Vogt et al. (2015)
Neighborhood intrastate conflicts Davies et al. (2024) Counterbalancing Pilster and Böhmelt (2011)
Democracy score Marshall and Gurr (2020) Military Exp. Barnum et al. (2025)
Coup attempt: Dummy Powell and Thyne (2011) Oil production Ross and Mahdavi (2015)
Coup attempt: Years since Powell and Thyne (2011) Gas production Ross and Mahdavi (2015)
Human rights Fariss (2019) Election year Hyde and Marinov (2012)
State capacity Hanson and Sigman (2021) Leader duration Mattes et al. (2016)
Bureaucratic capacity Coppedge et al. (2023) Regime duration Mattes et al. (2016)
Fiscal capacity Coppedge et al. (2023) Neighborhood unsucc. coups Powell and Thyne (2011)
Territorial control Coppedge et al. (2023) Neighborhood succ. coups Powell and Thyne (2011)
CSO entry and exit Coppedge et al. (2023) Int. rivalry: Dummy Thompson and Dreyer (2011)
CSO repression Coppedge et al. (2023) Int. rivalry: Count Thompson and Dreyer (2011)
CSO participatory env. Coppedge et al. (2023) MID: Dummy Palmer et al. (2022)
CSO anti-system mov. Coppedge et al. (2023) MID: Years since Palmer et al. (2022)
CSO strength Coppedge et al. (2023) MID: Count Palmer et al. (2022)
Regime change Mattes et al. (2016) Protest Chenoweth et al. (2014)
Urban population World Bank (2021) Neighborhood protest Chenoweth et al. (2014)
Economic growth World Bank (2021) GDP p.c. World Bank (2021)
Personalisation Geddes et al. (2018) Population World Bank (2021)
Ideological legitimisation Coppedge et al. (2023) Affiliated forces De Bruin (2021)
Neighborhood: Secret Police Choulis et al. (2024) Counterweight De Bruin (2021)
Appendix B Main Results
Table A.2 presents the full results for the four models underlying the marginal effects illustrated in Figure 1 of the main paper.
Table A.2: Full results table underlying figure 1. DV: Secret Police. Standard errors are reported in parentheses below the average marginal effects.
Variable LASSO Stepwise
Logit Cloglog Logit Cloglog
Intercept -8.7775 -8.4632 -11.2836 -10.9846